lime (++)*2 10 lim (cosk) 3/2 xoo use IBP to integrate x Sin (5x) dx x sin (25x) dx
Answer #1

Let

y=\lim_{x\rightarrow\infty}\left(1+\dfrac{1}x\right)^{x^2}

Taking log on both sides, we have

\log y= \lim_{x\rightarrow\infty}\log \left(1+\dfrac{1}x\right)^{x^2}

0041 + 1) 501,= f 801 =

\implies \log y= \lim_{x\rightarrow\infty} \dfrac{\log \left(1+\dfrac{1}x\right)}{1/x^2}\qquad (0/0 form)

Apply L' Hopital's Rule, we have

\implies \log y= \lim_{x\rightarrow\infty} \dfrac{ \left(0-\dfrac{1}{x^2}\right)}{(1+1/x)(-2/x^3)}

\implies \log y= \lim_{x\rightarrow\infty} \dfrac{ \left(x^2\right)}{2(x+1)}

Again apply L'Hopital's Rule, se have

\implies \log y= \lim_{x\rightarrow\infty} \dfrac{ \left(2x\right)}{2}

\implies \log y= \infty

\implies y=e^{ \infty}=\infty

Therefore, we have

y=\lim_{x\rightarrow\infty}\left(1+\dfrac{1}x\right)^{x^2}=\infty

10)

Let

y=\lim_{x\rightarrow0}\left(\cos x\right)^{3/x^2}

Taking log on both sides, we have

\log y=\log \lim_{x\rightarrow0}\left(\cos x\right)^{3/x^2}

\implies \log y=\lim_{x\rightarrow0}\log \left(\cos x\right)^{3/x^2}

\implies \log y=\lim_{x\rightarrow0}(3/x^2)\log \left(\cos x\right)

\implies \log y=3\lim_{x\rightarrow0}\dfrac{\log \left(\cos x\right)}{x^2}

\implies \log y=3\lim_{x\rightarrow 0}\dfrac{\log \left(\cos x\right)}{x^2}\qquad (0/0 form)

Apply L'Hopital's rule, we have

\implies \log y=3\lim_{x\rightarrow 0}\dfrac{\sin x}{\cos x (2x)}

\implies \log y=(3/2)\lim_{x\rightarrow 0}\dfrac{\sin x}x\dfrac{1}{\cos x }

\implies \log y=(3/2)(1)\dfrac{1}{\cos 0 }=(3/2)(1)(1)=3/2

\implies y=e^{3/2}

According to Chegg guidilines we cannot give the answer of all questions.

Please comment if you have any doubt or answer is mismatch.

Thank you.....